A bookshop owner purchased 100notebook for 15 each. However 5notebook had to be thrown away as a few pages were torn from them . Remaining books were sold at 18 each. Find loss or gain percent
.

Grade:8

1 Answers

RAJ KUMAR
6 years ago
total cost price=15×100
5 books thrown away..so 95 books left to be sold
total selling price=18×95=1710
%gain=((1710-1500)/1500)×100=14%